OpenStudy (anonymous):

I HAVE JUST A FEW MORE QUESTION PLEASE HELP???Solve the system of equations: what is the y-value in the solution? 3x - 5y = -16 and 2x + 5y = 31 5 3 2 4

OpenStudy (anonymous):

3x-5y=-16 2x+5y=31 So 3x=5y-16\[x=\frac{ 5y }{ 3 }-\frac{ 16 }{ 3 }\]Then\[2(\frac{ 5y }{ 3 }-\frac{ 16 }{ 3 })+5y=31\]\[\frac{ 10y }{ 3 }-\frac{ 32 }{ 3 }+\frac{ 15y }{ 3 }=31\]\[\frac{ 25y }{ 3 }=\frac{ 93 }{ 3 }+\frac{ 32 }{ 3 }\]\[\frac{ 25y }{ 3 }=\frac{ 125 }{ 3 }\]\[25y=125\]\[y\]

OpenStudy (anonymous):

y=5

OpenStudy (anonymous):

thsnk you